# Quantum-Corrected Thermodynamics of Conformal Weyl Gravity Black Holes: GUP Effects and Phase Transitions

### v1

@@
Abstract
We investigate the thermodynamic properties of black holes in Conformal Weyl Gravity (CWG) using the Mannheim-Kazanas solution, with particular emphasis on quantum corrections that become significant near the Planck scale. Our analysis employs the Hamilton-Jacobi tunneling formalism to derive the Hawking temperature, revealing explicit contributions from the conformal parameters $\beta$, $\gamma$, and $k$ that lead to substantial deviations from Schwarzschild black hole behavior. We incorporate quantum gravitational effects through the Generalized Uncertainty Principle, demonstrating systematic suppression of thermal radiation in the near-Planckian regime. Using an exponentially corrected entropy model, we compute the complete spectrum of QC thermodynamic potentials, including internal energy, pressure, heat capacity, and free energies. Our heat capacity analysis reveals critical phase transitions separating thermodynamically stable and unstable regions, with the scale-dependent parameter $\gamma$ playing a crucial role in controlling phase structure. The Joule-Thomson expansion analysis shows distinct cooling and heating regimes with inversion points that shift systematically with CWG parameters, capturing QC phase transitions absent in general relativity. We also examine gravitational redshift in CWG geometry, finding complex radial dependence that could provide observable signatures for distinguishing conformal gravity from Einstein’s theory. Our results demonstrate that CWG offers a consistent framework for studying black hole thermodynamics beyond general relativity, with quantum corrections revealing rich phase structures and potential observational consequences.
Black hole; Joule-Thomson Expansion; Quantum Correction; Redshift; Radiation; Generalized Uncertainty Principle; Conformal Weyl Gravity.
I Introduction
@@
CWG stands out prominently among alternative gravity theories due to its elegant local conformal invariance and sophisticated higher-derivative mathematical structure [3, 4, 5, 6, 7]. Unlike conventional GR, which constructs its gravitational action from the Ricci scalar $R$, CWG employs the square of the conformal Weyl tensor $C_{\mu\nu\rho\sigma}C^{\mu\nu\rho\sigma}$ as its fundamental building block, thereby introducing fourth-order field equations that admit significantly richer vacuum structures and more complex dynamical behaviors. The Mannheim-Kazanas solution, derived within this theoretical framework, represents a notable departure from classical Schwarzschild geometry through the inclusion of additional linear $\gamma r$ and quadratic $kr^{2}$ potential terms that emerge naturally from higher-order field equations [3, 8]. These non-Einsteinian contributions offer compelling explanations for various astrophysical observations, including galactic rotation curves and large-scale structure formation, without requiring the introduction of exotic dark matter components [9]. Furthermore, the presence of local conformal symmetry in CWE does not tolerate any dimensionful parameter (and hence no rigid scale) at sufficiently high energy scales. So, the Newton constant with dimension $[m^{-2}]$-which causes the irremediable catastrophic infinities at loop-level and hence leads to perturbatively non-renormalizable quantum gravity-comes into existence only in the classical vacuum of CWE where the conformal symmetry is broken.
@@
The incorporation of quantum corrections into BH thermodynamics represents another crucial frontier in our understanding of gravitational physics near the Planck scale, where classical descriptions inevitably break down and quantum effects become dominant [17, 18]. The Generalized Uncertainty Principle (GUP) [19, 20, 21, 22, 23, 24, 25], which emerges naturally from various approaches to quantum gravity including string theory, loop quantum gravity, and non-commutative geometry, provides a minimal yet powerful framework for capturing leading-order quantum gravitational effects through the introduction of a fundamental minimal length scale [26, 27, 28, 29, 30, 31]. This principle modifies the standard Heisenberg uncertainty relations and leads to systematic corrections in BH thermal properties, particularly affecting temperature, entropy, and heat capacity in ways that become increasingly important as BH masses approach Planck-scale values.
@@
The Joule-Thomson expansion (JTE) represents a particularly sensitive probe of BH thermodynamics, providing detailed information about phase transitions and thermal stability properties that are invisible to other thermodynamic measures [34, 35, 36, 37]. In the extended phase space formalism, where the cosmological constant is treated as a thermodynamic pressure, JTE analysis reveals rich phase structures including cooling and heating regimes separated by critical inversion points. These phenomena become even more complex in CWG due to the additional parameter space introduced by the conformal corrections. Gravitational redshift measurements offer another powerful avenue for testing alternative gravity theories, as they provide direct observational access to the spacetime metric structure near massive compact objects [38, 39]. The modification of redshift patterns in CWG compared to GR predictions could potentially be detected through high-precision spectroscopic observations of radiation emanating from the vicinity of supermassive BHs or neutron stars.
Our primary motivation for this comprehensive investigation stems from the recognition that understanding BH physics in alternative gravity theories, particularly when augmented with quantum corrections, represents a crucial step toward developing a complete theory of quantum gravity. CWG provides an ideal theoretical laboratory for this investigation due to its mathematical elegance, observational relevance, and rich phenomenological structure. Using the Mannheim-Kazanas metric as a gravitational background, we systematically investigate the thermodynamic behavior of BHs within CWG, with particular focus on quantum corrections that become significant near the Planck scale. We first revisit the semiclassical Hawking radiation through the Hamilton-Jacobi tunneling method [40, 41] and then incorporate corrections from the GUP, which is widely regarded as a minimal extension capturing leading-order quantum gravity effects [42, 43]. These corrections impact key thermodynamic quantities such as temperature, entropy, and heat capacity, opening possibilities for rich phase structures, including modifications to the JTE and redshift phenomena.
@@
The paper is organized as follows: In Section II, we provide a comprehensive review of the Mannheim-Kazanas BH geometry within the CWG framework, establishing the mathematical foundation for our subsequent analysis. Section III presents our evaluation of Hawking radiation using the Hamilton-Jacobi formalism for tunneling particles, revealing the thermal spectrum modifications introduced by conformal corrections. Section IV is devoted to the systematic calculation of quantum corrections arising from the GUP and their impact on thermal radiation properties. In Section V, we conduct a detailed study of quantum-corrected (QC) thermodynamics for CWGBHs, including the computation of all relevant thermodynamic potentials and their phase structure analysis. Section VI focuses on the QC JTE in CWGBHs, investigating isenthalpic processes and thermal stability properties. Section VII analyzes gravitational redshift phenomena in CWGBH geometry, exploring potential observational signatures of conformal corrections. Finally, Section VIII concludes our investigation by summarizing the main results and outlining prospective research directions that emerge from our findings.
II Review Of CWGBH Spacetime: Analysis of the Mannheim-Kazanas Solution
@@
$S=-\alpha_{g}\int d^{4}x\sqrt{-g}\,C_{\mu\nu\rho\sigma}C^{\mu\nu\rho\sigma},$ (1)
where $\alpha_{g}$ represents a dimensionless coupling constant that characterizes the strength of the conformal gravitational interaction. The conformal Weyl tensor in four spacetime dimensions, originally formulated by Weyl and later developed by Bach [52, 53], is explicitly expressed as
@@
where $R_{\mu\nu\rho\sigma}$ and $R_{\mu\nu}$ denote the Riemann and Ricci tensors respectively, while $R$ represents the scalar curvature. The conformally invariant Weyl tensor inherits the same fundamental symmetry properties as the ordinary Riemann tensor, satisfying the essential relations
$C_{\mu\nu\rho\sigma}=C_{[\mu\nu][\rho\sigma]}=C_{\rho\sigma\mu\nu},\quad C_{\mu[\nu\rho\sigma]}=0,\quad C^{\mu}{{}_{\nu\mu\sigma}}=0.$ (3)
It is noteworthy that this tensor identically vanishes in 2+12+12 + 1-dimensional spacetimes, where it is replaced by the Cotton tensor, highlighting the special role of four-dimensional geometry in conformal gravity theories [54].
The explicit form of the Weyl square term appearing in the action in Eq.(1) can be decomposed as [^1] By utilizing the topological Gauss-Bonnet invariant in conjunction with Eq.(5), the primary action in Eq.(1) admits an alternative representation: $S=-2\alpha_{g}\int d^{4}x\sqrt{-g}\,(R^{2}_{\mu\nu}-\frac{1}{3}R^{2}).$ (4)
@@
A crucial distinguishing feature of CWG compared to Einstein-Hilbert gravity lies in its respect for local conformal symmetry. The theory remains invariant under pointwise rescalings of the metric tensor according to $g_{\mu\nu}\rightarrow\Omega^{2}(x)g_{\mu\nu}$, where $\Omega(x)$ represents an arbitrary point-dependent positive function [^2] Under such local conformal transformations, the Weyl tensor transforms as $C_{\mu\nu\rho\sigma}\rightarrow C_{\mu\nu\rho\sigma}^{{}^{\prime}}=\Omega^{2}(x)C_{\mu\nu\rho\sigma}$, preserving the form of the action.. This enhanced symmetry structure has profound implications for the resulting field equations and vacuum solutions [3].
@@
$\frac{1}{\sqrt{-g}}\,\frac{\delta S}{\delta g_{\mu\nu}}=-2\alpha_{g}B^{\mu\nu}=-\frac{1}{2}T^{\mu\nu},$ (6)
@@
A fundamental observation is that in vacuum regions where the Ricci tensor vanishes, the Bach tensor $B^{\mu\nu}$ also vanishes identically. Consequently, all vacuum solutions of Einstein’s field equations automatically satisfy the CWG field equations, though the converse is not generally true, indicating that CWG admits a broader class of vacuum configurations than GR [3] [^4] This can be easily seen as one recast Eq.(7) into the form with Schouten tensor ${\cal P}_{\mu\nu}$ as $B_{\mu\nu}=2\nabla^{\alpha}\nabla_{[\alpha}{\cal P}_{\mu]\nu}-{\cal P}^{\alpha\beta}C_{\mu\alpha\nu\beta},$ (8) with ${\cal P}_{\mu\nu}=(1/2)(R_{\mu\nu}-(1/6)g_{\mu\nu}R)$ [74]. Since ${\cal P}_{\mu\nu}=(1/6)\Lambda g_{\mu\nu}$ in Einstein spaces (with $R_{\mu\nu}=\Lambda g_{\mu\nu}$ and $R=4\Lambda$), all the solutions of Einstein spaces are also those of CWE albeit its invalid reverse. For an informative study, see also [67].
@@
The radial structure of the metric function reveals distinct physical regimes: the term $-\beta(2-3\beta\gamma)/r$ dominates the Newtonian gravitational potential at short distances, the linear term $\gamma r$ becomes significant at intermediate scales, while the quadratic term $kr^{2}$ governs the behavior at cosmological distances. The zeros of $B(r)$ determine the locations of event horizons, and depending on the relative magnitudes and signs of the parameters $\beta$, $\gamma$, and $k$, the solution can exhibit multiple real, positive roots corresponding to distinct BH horizons.
From a geometric perspective, the Mannheim-Kazanas or CWGBH metric represents a non-trivial vacuum solution of the fourth-order Bach equations, which replace Einstein’s second-order field equations in CWG. These higher-order equations admit more general vacuum configurations due to their enhanced derivative structure and enlarged symmetry content. The associated curvature invariants reveal the presence of curvature singularities at $r=0$, maintaining consistency with classical BH behavior, while the causal structure defined by the horizon locations provides a rich framework for studying thermodynamic properties.
III Hawking Radiation of CWGBHs via Hamilton-Jacobi Formalism of Tunneling Particles
The investigation of quantum tunneling processes for scalar particles in curved spacetime geometries provides a powerful framework for understanding thermal radiation from BHs. The relativistic Hamilton-Jacobi equation offers an elegant and computationally tractable approach to analyze such phenomena within the context of CWG. Given the spacetime geometry characterized by the Mannheim-Kazanas solution in Eq.(10), we examine the behavior of a scalar field described by an action $S$ that satisfies the fundamental constraint [75]
@@ III Hawking Radiation of CWGBHs via Hamilton-Jacobi Formalism of Tunneling Particles
$g^{tt}=-\frac{1}{B(r)},\quad g^{rr}=B(r),\quad g^{\theta\theta}=\frac{1}{r^{2}},\quad g^{\phi\phi}=\frac{1}{r^{2}\sin^{2}\theta},$ (12)
where $B(r)$ represents the radial metric function derived from the CWG field equations. Substituting these metric components into the Hamilton-Jacobi equation yields the explicit form
$-\frac{(\partial_{t}S)^{2}}{B(r)}+B(r)(\partial_{r}S)^{2}+\frac{(\partial_{\theta}S)^{2}}{r^{2}}+\frac{(\partial_{\phi}S)^{2}}{r^{2}\sin^{2}\theta}+m^{2}=0.$ (13)
@@
$-\frac{E^{2}}{B(r)}+B(r)\left(\frac{dW}{dr}\right)^{2}+m^{2}=0,$ (16)
@@
This integral exhibits a characteristic logarithmic divergence at the horizon, which gives rise to an imaginary contribution to the action. This imaginary part is crucial for the tunneling interpretation and is given by
@@
$T_{H}=\frac{B^{\prime}(r_{h})}{4\pi}=\frac{3\beta^{2}\gamma}{4\pi r_{h}^{2}}+\frac{\beta}{2\pi r_{h}^{2}}+\frac{\gamma}{4\pi}+\frac{kr_{h}}{2\pi}.$ (26)

### v2

@@
Abstract. We investigate the thermodynamic properties of black holes in Conformal Weyl Gravity (CWG) using the Mannheim-Kazanas solution, with particular emphasis on quantum corrections that become significant near the Planck scale. Our analysis employs the Hamilton-Jacobi tunneling formalism to derive the Hawking temperature, revealing explicit contributions from the conformal parameters $\beta$, $\gamma$, and $k$ that lead to substantial deviations from Schwarzschild black hole behavior. We incorporate quantum gravitational effects through the Generalized Uncertainty Principle, demonstrating systematic suppression of thermal radiation in the near-Planckian regime. Using an exponentially corrected entropy model, we compute the complete spectrum of QC thermodynamic potentials, including internal energy, pressure, heat capacity, and free energies. Our heat capacity analysis shows divergence behavior that separates stable and unstable regions, indicating possible thermodynamic transitions controlled by the scale-dependent parameter $\gamma$. The Joule-Thomson expansion analysis shows distinct cooling and heating regimes with inversion points that shift systematically with CWG parameters, capturing QC phase transitions absent in general relativity. We also examine gravitational redshift in CWG geometry, finding complex radial dependence that highlights modifications compared to the Schwarzschild case, although redshift alone cannot observationally distinguish CWG from Einstein’s theory. Our results demonstrate that CWG offers a consistent framework for studying black hole thermodynamics beyond general relativity, with quantum corrections modifying phase structures in the near-Planckian regime, though these effects are not expected to yield direct observational consequences.
Keywords: Black hole; Joule-Thomson Expansion; Quantum Correction; Redshift; Radiation; Generalized Uncertainty Principle; Conformal Weyl Gravity.
1 Introduction
@@
CWG stands out prominently among alternative gravity theories due to its elegant local conformal invariance and sophisticated higher-derivative mathematical structure [85, 84, 80, 79, 83]. Unlike conventional GR, which constructs its gravitational action from the Ricci scalar $R$, CWG employs the square of the conformal Weyl tensor $C_{\mu\nu\rho\sigma}C^{\mu\nu\rho\sigma}$ as its fundamental building block, thereby introducing fourth-order field equations that admit significantly richer vacuum structures and more complex dynamical behaviors. The Mannheim-Kazanas (MK) solution, derived within this theoretical framework, represents a notable departure from classical Schwarzschild geometry through the inclusion of additional linear $\gamma r$ and quadratic $kr^{2}$ potential terms that emerge naturally from higher-order field equations [85, 82]. These non-Einsteinian contributions offer compelling explanations for various astrophysical observations, including galactic rotation curves and large-scale structure formation, without requiring the introduction of exotic dark matter components [87]. Furthermore, the presence of local conformal symmetry in CWE does not tolerate any dimensionful parameter (and hence no rigid scale) at sufficiently high energy scales. So, the Newton constant with dimension $m^{-2}]-$which causes the irremediable catastrophic infinities at loop-level and hence leads to perturbatively non-renormalizable quantum gravity-comes into existence only in the classical vacuum of CWE where the conformal symmetry is broken.
@@
The incorporation of quantum corrections into BH thermodynamics represents another crucial frontier in our understanding of gravitational physics near the Planck scale, where classical descriptions inevitably break down and quantum effects become dominant [30, 40]. The Generalized Uncertainty Principle (GUP) [50, 114, 109, 131, 145, 64, 144, 115, 2, 135], which emerges naturally from various approaches to quantum gravity including string theory, loop quantum gravity, and non-commutative geometry, provides a minimal yet powerful framework for capturing leading-order quantum gravitational effects through the introduction of a fundamental minimal length scale [105, 48, 95, 49, 96, 51, 133]. This principle modifies the standard Heisenberg uncertainty relations and leads to systematic corrections in BH thermal properties, particularly affecting temperature, entropy, and heat capacity in ways that become increasingly important as BH masses approach Planck-scale values.
@@
The Joule-Thomson expansion (JTE) represents a particularly sensitive probe of BH thermodynamics, providing detailed information about phase transitions and thermal stability properties that are invisible to other thermodynamic measures [88, 99, 36, 100, 15]. In the extended phase space formalism, where the cosmological constant is treated as a thermodynamic pressure, JTE analysis reveals rich phase structures including cooling and heating regimes separated by critical inversion points. These phenomena become even more complex in CWG due to the additional parameter space introduced by the conformal corrections. Gravitational redshift measurements offer another powerful avenue for testing alternative gravity theories, as they provide direct observational access to the spacetime metric structure near massive compact objects [57, 154]. The modification of redshift patterns in CWG compared to GR predictions could potentially be detected through high-precision spectroscopic observations of radiation emanating from the vicinity of supermassive BHs or neutron stars.
Our primary motivation for this comprehensive investigation stems from the recognition that understanding BH physics in alternative gravity theories, particularly when augmented with quantum corrections, represents a crucial step toward developing a complete theory of quantum gravity. CWG provides an ideal theoretical laboratory for this investigation due to its mathematical elegance, observational relevance, and rich phenomenological structure. Using the MK metric as a gravitational background, we systematically investigate the thermodynamic behavior of BHs within CWG, with particular focus on quantum corrections that become significant near the Planck scale. We first revisit the semiclassical Hawking radiation through the Hamilton-Jacobi tunneling method [103, 134, 127] and then incorporate corrections from the GUP, which is widely regarded as a minimal extension capturing leading-order quantum gravity effects [3, 97]. These corrections impact key thermodynamic quantities such as temperature, entropy, and heat capacity, opening possibilities for rich phase structures, including modifications to the JTE and redshift phenomena.
@@
The paper is organized as follows: In Section 2, we provide a comprehensive review of the MK BH geometry within the CWG framework, establishing the mathematical foundation for our subsequent analysis. Section 3 presents our evaluation of Hawking radiation using the Hamilton-Jacobi formalism for tunneling particles, revealing the thermal spectrum modifications introduced by conformal corrections. Section 4 is devoted to the systematic calculation of quantum corrections arising from the GUP and their impact on thermal radiation properties. In Section 5, we conduct a detailed study of quantum-corrected (QC) thermodynamics for CWGBHs, including the computation of all relevant thermodynamic potentials and their phase structure analysis. Section 6 focuses on the QC JTE in CWGBHs, investigating isenthalpic processes and thermal stability properties. Section 7 analyzes gravitational redshift phenomena in CWGBH geometry, exploring potential observational signatures of conformal corrections. Finally, Section 8 concludes our investigation by summarizing the main results and outlining prospective research directions that emerge from our findings.
2 Review Of CWGBH Spacetime: Analysis of MK Solution
@@
$\mathcal{S}=-\alpha_{g}\int d^{4}x\sqrt{-g}\,C_{\mu\nu\rho\sigma}C^{\mu\nu\rho\sigma},$ (1)
where $\alpha_{g}$ represents a dimensionless coupling constant that characterizes the strength of the conformal gravitational interaction. This action is conformally invariant specifically in four spacetime dimensions, since the conformal weight of $C_{\mu\nu\rho\sigma}C^{\mu\nu\rho\sigma}$ (weight 4-4) is exactly compensated by $\sqrt{-g}\to\Omega^{4}\sqrt{-g}$ only for $D=4$; in general $D$ dimensions, $\sqrt{-g}\to\Omega^{D}\sqrt{-g}$ while $C^{2}$ transforms with weight 4-4 regardless of $D$, so the cancellation fails for $D\neq 4$ [148, 45]. The conformal Weyl tensor in four spacetime dimensions, originally formulated by Weyl and later developed by Bach [152, 16], is explicitly expressed as
@@
where $R_{\mu\nu\rho\sigma}$ and $R_{\mu\nu}$ denote the Riemann and Ricci tensors respectively, while $R$ represents the scalar curvature. The conformally covariant Weyl tensor inherits the same fundamental symmetry properties as the ordinary Riemann tensor, satisfying the essential relations
$C_{\mu\nu\rho\sigma}=C_{[\mu\nu][\rho\sigma]}=C_{\rho\sigma\mu\nu},$ (3)
$C_{\mu[\nu\rho\sigma]}=0,\qquad C^{\mu}{{}_{\nu\mu\sigma}}=0.$ (4)
It is noteworthy that this tensor identically vanishes in $2+1$-dimensional spacetimes, where it is replaced by the Cotton tensor, highlighting the special role of four-dimensional geometry in conformal gravity theories [112].
The explicit form of the Weyl square term appearing in the action in Eq. (1) can be decomposed as [^4] By utilizing the topological Gauss-Bonnet invariant in conjunction with Eq. (5), the primary action in Eq. (1) admits an alternative representation: $\mathcal{S}=-2\alpha_{g}\int d^{4}x\sqrt{-g}\,(R^{2}_{\mu\nu}-\tfrac{1}{3}R^{2}).$
@@
A crucial distinguishing feature of CWG compared to Einstein-Hilbert gravity lies in its respect for local conformal symmetry. The theory remains invariant under pointwise rescalings of the metric tensor according to $g_{\mu\nu}\rightarrow\Omega^{2}(x)g_{\mu\nu}$, where $\Omega(x)$ represents an arbitrary point-dependent positive function [^5] Under such local conformal transformations, the Weyl tensor transforms as $C_{\mu\nu\rho\sigma}\rightarrow C_{\mu\nu\rho\sigma}^{{}^{\prime}}=\Omega^{2}(x)C_{\mu\nu\rho\sigma}$, which constitutes conformal covariance (weight +2+2), not invariance; only the mixed-index form $C^{\mu}{}_{\nu\rho\sigma}$ is conformally invariant. The Weyl-squared scalar $C_{\mu\nu\rho\sigma}C^{\mu\nu\rho\sigma}$ transforms with weight $\Omega^{-4}$, which is compensated by $\sqrt{-g}\to\Omega^{4}\sqrt{-g}$ in four dimensions, preserving the form of the action.. This enhanced symmetry structure has profound implications for the resulting field equations and vacuum solutions [85].
@@
$\frac{1}{\sqrt{-g}}\,\frac{\delta\mathcal{S}}{\delta g_{\mu\nu}}=-2\alpha_{g}B^{\mu\nu}=-\frac{1}{2}T^{\mu\nu},$ (6)
@@
A fundamental observation is that in vacuum regions where the Ricci tensor vanishes, the Bach tensor $B^{\mu\nu}$ also vanishes identically. Consequently, all vacuum solutions of Einstein’s field equations automatically satisfy the CWG field equations, though the converse is not generally true, indicating that CWG admits a broader class of vacuum configurations than GR [85] [^7] This can be easily seen as one recasts Eq. (7) into the form with Schouten tensor $\mathcal{P}_{\mu\nu}$ as $B_{\mu\nu}=2\nabla^{\alpha}\nabla_{[\alpha}\mathcal{P}_{\mu]\nu}-\mathcal{P}^{\alpha\beta}C_{\mu\alpha\nu\beta}$, with $\mathcal{P}_{\mu\nu}=(1/2)(R_{\mu\nu}-(1/6)g_{\mu\nu}R)$ [23]. Since $\mathcal{P}_{\mu\nu}=(1/6)\Lambda g_{\mu\nu}$ in Einstein spaces (with $R_{\mu\nu}=\Lambda g_{\mu\nu}$ and $R=4\Lambda$), all the solutions of Einstein spaces are also those of CWE albeit its invalid reverse. For an informative study, see also [63].
@@
The radial structure of the metric function reveals distinct physical regimes: the term $\beta(2-3\beta\gamma)/r$ dominates the Newtonian gravitational potential at short distances, the linear term $\gamma r$ becomes significant at intermediate scales, while the quadratic term $kr^{2}$ governs the behavior at cosmological distances. The zeros of $B(r)$ determine the locations of event horizons, and depending on the relative magnitudes and signs of the parameters $\beta$, $\gamma$, and $k$, the solution can exhibit multiple real, positive roots corresponding to distinct BH horizons.
From a geometric perspective, the MK or CWGBH metric represents a non-trivial vacuum solution of the fourth-order Bach equations, which replace Einstein’s second-order field equations in CWG. These higher-order equations admit more general vacuum configurations due to their enhanced derivative structure and enlarged symmetry content. The associated curvature invariants reveal the presence of curvature singularities at $r=0$, maintaining consistency with classical BH behavior, while the causal structure defined by the horizon locations provides a rich framework for studying thermodynamic properties.
3 Hawking Radiation of CWGBHs via Hamilton-Jacobi Formalism of Tunneling Particles
The investigation of quantum tunneling processes for scalar particles in curved spacetime geometries provides a powerful framework for understanding thermal radiation from BHs. The relativistic Hamilton-Jacobi equation offers an elegant and computationally tractable approach to analyze such phenomena within the context of CWG. Given the spacetime geometry characterized by the MK solution in Eq. (9), we examine the behavior of a scalar field described by an action $S$ that satisfies the fundamental constraint [123]
@@ 3 Hawking Radiation of CWGBHs via Hamilton-Jacobi Formalism of Tunneling Particles
$g^{tt}$ $=-B(r),\quad g^{rr}=\frac{1}{B(r)},$
$g^{\theta\theta}$ $=\frac{1}{r^{2}},\quad g^{\phi\phi}=\frac{1}{r^{2}\sin^{2}\theta},$ (11)
where, as mentioned above, $B(r)$ represents the radial metric function derived from the CWG field equations. Substituting these metric components into the Hamilton-Jacobi equation yields the explicit form
$\frac{(\partial_{t}S)^{2}}{B(r)}+B(r)(\partial_{r}S)^{2}+\frac{(\partial_{\theta}S)^{2}}{r^{2}}+\frac{(\partial_{\phi}S)^{2}}{r^{2}\sin^{2}\theta}+m^{2}=0.$ (12)
@@
$\frac{E^{2}}{B(r)}+B(r)\left(\frac{dW}{dr}\right)^{2}+m^{2}=0,$ (15)
@@
This integral exhibits a characteristic logarithmic divergence at the horizon, which gives rise to an imaginary contribution to the action. This imaginary part is crucial for the tunneling interpretation and is given by [132]
@@
$T_{H}=\frac{B^{\prime}(r_{h})}{4\pi}=-\frac{3\beta^{2}\gamma}{4\pi r_{h}^{2}}+\frac{\beta}{2\pi r_{h}^{2}}+\frac{\gamma}{4\pi}+\frac{kr_{h}}{2\pi}.$ (25)